if 2nC3 : nC2= 12 : 1 then what will be the value of n? pls give answer with solution

Grade:12th pass

1 Answers

imran
7 years ago
2nC3/nC2  =  12/1
((2n)!/3!*(2n-3)! )  / n!*2!(n-2)! =12
(2(2n-1)(2n-2) )/3(n-1)  =12
4(2n-1) =36
2n =10
 
There fore n=5